# DeFT: Decoding with Flash Tree-Attention for Efficient Tree-structured LLM Inference

### v3

@@ 1 Introduction
Table 1: Comparison of efficiency in sequence-based (CoT (Wei et al., 2022)) and tree-based (ToT (Yao et al., 2023)) decoding for a reasoning task. The task is sorting $128$ numbers from Besta et al. (2023). The total generated tokens of CoT is only $525$ while $38{,}315$ in ToT, resulting in inefficiency in end-to-end latency (second) and IO (TB). IO mainly consists of two parts as follows. (i) KV cache: IO-KV; (ii) Partial results during attention calculation like $QK^{T}$ and softmax: IO-PA; Baselines: (i) Flash-Decoding (Dao et al., 2023); (ii) Tree Attention: tree attention in Medusa (Cai et al., 2024).
Metrics
@@
DeFT-Flatten(ours) + ToT 94.61 12.40 0
@@
  • We empirically verify its effectiveness on few-shot prompting, multi-step reasoning, and speculative-decoding tasks. DeFT-Flatten can achieve a walk-clock time speedup of 1.3$\times$ for few-shot prompting, 2.5$\times$ for speculative decoding, 1.1$\times$ for multi-step reasoning, due to an up to 3.82$\times$ faster attention calculation, with the baseline implementations (Dao et al., 2023; Cai et al., 2024; Zheng et al., 2023).
@@ 2 Related Work
Discussion on Concurrent Works. Some concurrent works (Ye et al., 2024; Juravsky et al., 2024; Athiwaratkun et al., 2024) also recognize the importance of IO during LLM inference. However, these works have at least one of these flaws: i) they (Ye et al., 2024; Juravsky et al., 2024; Athiwaratkun et al., 2024) cannot be easily extended to situations where the decoding tree has more than two levels—they target single-context batch sampling scenarios, a special case of general tree-based decoding with a system prompt as prefix and unique suffixes in the first depth; ii) they (Juravsky et al., 2024; Athiwaratkun et al., 2024) do not consider the inefficiency caused by the lengths of different nodes in the decoding tree. See the comparison of DeFT and concurrent works in Appendix A.3.
@@ 3.3 Prefix-aware and Balanced Tree-structured KV Cache Partitions
Figure 3: Comparison of QKV partitioning strategies during the QKV Preparation Phase and influence on memory access for different attention algorithms. The decoding example is shown on the left of Figure 2. Note that the partitioning is logically designed without incurring any data movement costs for QKV. The partitioning strategy will guide the loading of QKV during the subsequent Attention calculation phase, where each QKV group $G_{i}$ will be loaded into $SM_{i}$ on the GPU. The amount of IO between the GPU HBM and shared memory required by each group is highlighted in red rectangles. (Left) From top to bottom, the figure illustrates notations, Vanilla Tree Attention, and most importantly, the details of Flatten Tree KV Splitting (see Remark 3.1). (Right) Different partitioning strategies lead to varying memory access amounts and load-balancing levels. DeFT-Flatten is better than others due to prefix-aware and balanced partitions. Q-Guided Grouping (Dao et al., 2022; 2023; Zheng et al., 2023; Miao et al., 2023), introduces significant redundancy in KV cache access. For a summary of baselines and DeFT, see Table 2. More details on IO complexity are provided in Appendix A.5.
@@
Let’s begin with a decoding example using the tree-structured KV cache shown in Figure 2. If we group the entire tree-structured KV cache and queries into ${\bm{G}}_{0}$ without any partitions, we can refer to the Vanilla Tree Attention method illustrated in the left part of Figure 3. This method calculates attention for all queries simultaneously in a single Streaming Multiprocessor (SM), with the aid of a dense causal mask (DCM).
@@
Table 2: Comparison of QKV partitioning strategies for baselines (most of which are shown in Figure 3) and DeFT. For IO redundancy, significant issues are highlighted in red, while negligible ones are in blue. “Q” refers to queries, and “KV” refers to the KV cache. “DCM” stands for Dense Causal Mask (a matrix), and “BCM” refers to Bit Causal Mask (a set of 64-bit integers). “PA” represents partial results during attention calculations, including ${\bm{Q}}{\bm{K}}^{T}$, Softmax, etc. More $\star$ symbols indicate better-balanced workloads for QKV partitions. Details on IO complexity can be found in Appendix A.5. See Figure 11 for tree attention partitioning and memory access.
@@
As shown in the bottom left of Figure 3, there are three key components:
@@
Several existing attention algorithms are designed for speculative decoding, where attention is calculated for the entire tree-structured queries. However, these methods are not memory-efficient. For a detailed IO analysis, refer to Appendix A.5.
  • Tree Attention-Medusa (Cai et al., 2024). Based on Vanilla Tree Attention (shown on the left in Figure 3), this method uses PyTorch’s General Matrix Multiply (GEMM) to partition Q and KV tensors. It is memory-inefficient for two reasons: (1) it does not utilize Flash-Attention to reduce memory access during the computation of intermediate results (e.g., Softmax); (2) it introduces a dense causal mask, whose memory access is significant (see Figure 11).
  • Tree Attention-SpecInfer (Miao et al., 2023). This algorithm employs Q-Guided Grouping based on Vanilla Tree Attention and partitions the KV sequence through Flash-Decoding. It is memory-inefficient in redundantly loading the entire tree-structured KV cache for each query. See Figure 11.
@@ 4.1 Experimental Setup
Few-shot prompting APPS - - 400 iterations
Multi-step reasoning 4 tasks in GoT ToT-BFS Reconstruct from interaction records with GPT 3.5 in GoT End of task
@@
Table 5: Comparison of DeFT-Flatten and baselines in average end-to-end latency (in seconds) for tree-based decoding. Here, $b$ represents the tree width, and $t$ denotes the token tree size (i.e., the number of tree-structured queries). The fastest method is in bold, and the second fastest is underlined. Speedup over the best wall-clock time refers to the speedup of DeFT-Flatten compared to the best baseline (Radix Attention) in terms of end-to-end latency. $\star$ denotes out-of-memory (OOM) errors for the A100 80GB GPU. Speedup Upper-bound (no attention) refers to the maximum speedup we could achieve for the best wall-clock latency baseline (Radix Attention) if we exclude the attention computation. Attention Speedup over the best wall-clock indicates the attention speedup of DeFT-Flatten over the best wall-clock latency baseline (Radix Attention). For more details on attention speedup, see Table 16.
@@
Paged Radix Attention 12.37 14.08 16.54 104.79 69.61 11.25 17.03 64.57 86.12 145.88 263.76
DeFT-Flatten 9.98 10.99 12.48 94.67 66.95 10.90 16.10 44.94 50.48 65.44 104.65
Attention Speedup over the best wall-clock 1.73$\times$ 1.63$\times$ 1.70$\times$ 1.39$\times$ 1.15$\times$ 1.21$\times$ 1.34$\times$ 2.57$\times$ 3.00×3.$0$ 3.64×3.64$\times$3.64 × 3.82×3.82$\times$3.82 ×
Speedup over the best wall-clock 1.24$\times$ 1.28$\times$ 1.33$\times$ 1.10$\times$ 1.03$\times$ 1.03$\times$ 1.05$\times$ 1.43$\times$ 1.70×1.70$\times$1.70 × $2\times2$ 2.52×2.52$\times$2.52 ×
Upper-bound(no attention) 1.71$\times$ 2.08$\times$ 2.51$\times$ 1.96$\times$ 1.82$\times$ 1.70$\times$ 1.76$\times$ 2.01$\times$ 2.72$\times$ 3.99$\times$ 5.12$\times$
@@ 4.2 Analysis of Memory Management and Bottleneck
Figure 4: Latency breakdown for speculative decoding with a token tree of 32 queries, whose tree topology is from Medusa (Cai et al., 2024). U means unpaged memory management.
@@
4.3 End-to-end Behaviors: Latency and IOs.
We evaluate DeFT’s performance on various tree-based decoding tasks by measuring end-to-end latency (Table 5). See attention latency (Table 16), IO (Table 17), and inference accuracy (Table 15) in Appendix A.7. This assessment demonstrates DeFT’s optimization of tree attention and its acceleration of wall-clock time.
For few-shot prompting tasks, we used a prompt with 4k tokens and performed 400 decoding iterations, achieving a 1.33$\times$ end-to-end speedup thanks to 1.70$\times$ faster attention calculation and an approximately 90% reduction in IO.
For speculative decoding tasks, DeFT-Flatten achieved up to a 2.52$\times$ wall-clock time speedup due to up to a 3.82$\times$ attention speedup, as the entire token tree (all queries) can share IO of the long prefix.
[…]
@@ 4.4 Ablation Study
Table 6: [Different KV Splitting Strategies] Comparison of DeFT-Node, DeFT-Node-Chunk and DeFT-Flatten in average attention latency (second) with NVIDIA A100 (80GB) for Llama3-8B model(GQA). This table is supplementary to Table 16. The fastest method is in bold, and the second fastest is underlined. Radix Attention is the best baseline in end-to-end latency. See details of more baselines in Table 16.
@@
Paged Radix Attention 5.99 7.30 9.96 39.37 24.69 3.11 5.13 32.60 54.57 109.39 212.29
DeFT-Node 10.59 10.62 10.85 42.96 33.29 6.16 9.58 50.82 52.51 53.06 61.82
DeFT-Node-Chunk 8.52 9.69 13.45 49.63 36.37 4.77 7.40 22.45 46.40 77.72 143.65
DeFT-Flatten 3.47 4.07 5.87 28.41 21.45 2.57 3.83 12.68 18.18 29.97 55.58
@@
See Table 8. With a longer prompt, DeFT-Flatten shows a more pronounced speedup in the same model, since the attention overhead is proportional to the token count in the decoding tree, while the FFN overhead remains nearly constant for the same model.
@@
See Table 8. With the larger model, Codellama-34B, DeFT-Flatten achieves slightly reduced but significant (up to 1.78$\times$) end-to-end speedup. The performance reduction is attributed to a lower A/F-LR, as the FFN overhead is greater due to larger hidden dimensions.
@@
End-to-end 1.09$\times$ 1.37$\times$ 1.53$\times$ 1.67$\times$
Table 8: [Different Model Sizes] Comparison of wall-clock time speedup and Attention/FFN latency ratio (in short as A/F-LR) between DeFT and Radix Attention for Codellama-34B and Codellama-7B models. Radix Attention is the best baseline in end-to-end latency. $b$ represents the tree width, and $t$ denotes the token tree size. For multi-step reasoning, we test the task sorting whose prompt length is about 1k tokens.
@@
Wall Clock Time Speedup 7B 1.34$\times$ 1.09$\times$ 1.85$\times$
@@ 5 Conclusion
We propose DeFT-Flatten, a hardware-efficient attention algorithm optimized for tree-structured LLM inference. It effectively addresses memory access and GPU utilization bottlenecks by reusing shared prefixes’ KV cache and evenly distributing workload across partitions. DeFT-Flatten’s key strengths lie in its prefix-sharing awareness and load balancing, making it versatile for various tree-structured tasks. It also scales well with larger search spaces and multiple branches. Our results show that DeFT-Flatten achieves up to 2.5x/3.82x speedup in end-to-end and attention latency, outperforming baselines in tasks such as few-shot prompting, multi-step reasoning, and speculative decoding. Our ablation studies highlight that: (1) balanced partitioning is critical, (2) DeFT-Flatten delivers significant speedups across various LLM models and GPU architectures, and (3) DeFT-Flatten provides even greater speedups with larger token sizes (e.g., longer prompts) and more branches in tree-structured requests.
@@ A.2 Discussion of Tree-based Decoding
The existing inference frameworks (Zheng et al., 2023; Gim et al., 2023) focused on tree-based decoding efficiency primarily aim to: (1) reduce memory footprints (Zheng et al., 2023) to enable larger batch sizes for higher throughput; (2) reuse the prompt cache (Gim et al., 2023) to avoid recomputation of the KV cache for faster time-to-first-token (TTFT). However, their designs do not specifically target reducing the wall-clock time of the entire decoding process. We observe that the tree-structured feature of LLM inference could provide us with some advantages to speed up the decoding itself.

### v4

@@ 1 Introduction
Table 1: Comparison of efficiency in sequence-based CoT (Wei et al., 2022) and tree-based ToT (Yao et al., 2023) decoding for a reasoning task. The task is sorting $128$ numbers from Besta et al. (2023). The total generated tokens of CoT is only $525$ while $38{,}315$ in ToT, resulting in inefficiency in end-to-end latency (second) and IO (TB). IO mainly consists of two parts as follows. (i) KV cache: IO-KV; (ii) Partial results during attention calculation like $QK^{T}$ and softmax: IO-PA; Baselines: (i) Flash-Decoding (Dao et al., 2023); (ii) Tree Attention: tree attention in Medusa (Cai et al., 2024).
@@
DeFT-Flatten(ours) + ToT 94.67 12.40 0
@@
  • We empirically verify its effectiveness on few-shot prompting, multi-step reasoning, and speculative-decoding tasks. DeFT-Flatten can achieve a decoding latency speedup of 1.3$\times$ for few-shot prompting, 2.2$\times$ for speculative decoding, 1.1$\times$ for multi-step reasoning, due to an up to 3.59$\times$ faster attention calculation, with the baseline implementations (Dao et al., 2023; Cai et al., 2024; Zheng et al., 2023).
@@ 2 Related Work
Discussion on Concurrent Works. Some concurrent works (Ye et al., 2024a; Juravsky et al., 2024; Athiwaratkun et al., 2024; Ye et al., 2024b; Zhu et al., 2024) also recognize the importance of IO during LLM inference. However, these works have at least one of these flaws: i) they (Ye et al., 2024a; Juravsky et al., 2024; Athiwaratkun et al., 2024; Ye et al., 2024b; Zhu et al., 2024) cannot be easily extended to situations where the decoding tree has more than two levels—they target single-context batch sampling scenarios, a special case of general tree-based decoding with a system prompt as prefix and unique suffixes in the first depth; ii) they (Juravsky et al., 2024; Athiwaratkun et al., 2024) do not consider the inefficiency caused by the lengths of different nodes in the decoding tree. See the comparison of DeFT and concurrent works in Appendix A.3.
@@ 3.3 Prefix-aware and Balanced Tree-structured KV Cache Partitions
Figure 3: Comparison of QKV partitioning strategies during the QKV Preparation Phase between DeFT-Node/Node-Chunk/Flatten and different attention algorithm baselines. Note that the partitioning is logically designed without incurring any data movement costs for QKV. The amount of IO between the GPU HBM and shared memory required by each group is highlighted in red rectangles. Part (a) illustrates the dataflow of a two-cascaded decoding tree example and three categories of QKV partitioning strategies: no partition(Vanilla Tree Attention), Q-Guided Grouping and KV-Guided Grouping. The partitioning strategy will guide the loading of QKV during the subsequent Attention calculation phase, where each QKV group $G_{i}$ will be loaded into $SM_{i}$ on the GPU. Part (b) shows the comparison of Q-Guided Grouping and KV-Guided Grouping, where the latter can be IO-aware of prefix KV cache $KV_{0}$ and only load it once. DeFT-Node-Chunk is a weak load-balancing improvement of DeFT-Node by splitting large nodes (e.g., $KV_{0}$) to chunks. Part (c) illustrates the details (discussed in Remark 3.1) of Flattened Tree KV Splitting in DeFT-Flatten for load-balanced partitions, including Depth-first Flatten strategy, Evenly block-wise strategy, and Bit mask. For a summary of baselines and DeFT, see Table 2. See analysis of tree-attention baselines (Cai et al., 2024; Miao et al., 2023) in Remark 3.2.
@@
Let’s begin with a decoding example using the tree-structured KV cache shown in Figure 2. If we group the entire tree-structured KV cache and queries into ${\bm{G}}_{0}$ without any partitions, we can refer to the Vanilla Tree Attention method illustrated in the part (a) of Figure 3. This method calculates attention for all queries simultaneously in a single Streaming Multiprocessor (SM), with the aid of a dense causal mask (DCM).
@@
Table 2: Comparison of QKV partitioning strategies for baselines (most of which are shown in Figure 3) and DeFT. For IO redundancy, significant issues are highlighted in red, while negligible ones are in blue. “Q” refers to queries, and “KV” refers to the KV cache. “DCM” stands for Dense Causal Mask (a matrix), and “BCM” refers to Bit Causal Mask (a set of 64-bit integers). “PA” represents partial results during attention calculations, including ${\bm{Q}}{\bm{K}}^{T}$, Softmax, etc. More $\star$ symbols indicate better-balanced workloads for QKV partitions. Details on IO complexity can be found in Appendix A.5.
@@
As shown in the part (c) of Figure 3, there are three key components:
@@
Existing attention algorithms are designed for speculative decoding, where attention is calculated for the entire tree-structured queries. However, these methods are not memory-efficient. For partition details, see Figure 11 in Appendix A.5.
  • Tree Attention-Medusa (Cai et al., 2024). Based on Vanilla Tree Attention (shown on the left in Figure 3), this method uses PyTorch’s General Matrix Multiply (GEMM) to partition Q and KV tensors. It is memory-inefficient for two reasons: (1) it does not utilize Flash-Attention to reduce memory access during the computation of intermediate results (e.g., Softmax); (2) it introduces a dense causal mask, whose memory access is significant.
  • Tree Attention-SpecInfer (Miao et al., 2023). This algorithm employs Q-Guided Grouping based on Vanilla Tree Attention and partitions the KV sequence through Flash-Decoding. It is memory-inefficient in redundantly loading the entire tree-structured KV cache for each query.
@@ 4.1 Experimental Setup
Few-shot prompting APPS - Pad the prompt to 4000 tokens 400 iterations
Multi-step reasoning 4 tasks in GoT ToT-BFS Reconstruct from interaction records with GPT 3.5 in GoT End of task($\sim$ 3500 iterations)
@@
Table 5: Comparison of DeFT-Flatten and baselines in average decoding latency (in seconds) for tree-based decoding. Here, $b$ represents the tree width, and $t$ denotes the token tree size (i.e., the number of tree-structured queries). The fastest method is in bold, and the second fastest is underlined. Radix Attention is the best baseline in decoding latency. $\star$ denotes out-of-memory (OOM) errors for the A100 80GB GPU. Speedup Upper-bound (no attention) refers to the maximum speedup we could achieve for Radix Attention if we exclude the attention computation and only run other components including MLP. For more details on attention speedup, see Table 16.
@@
Paged Radix Attention 12.37 14.08 16.54 104.79 69.61 11.25 17.03 54.66 69.75 108.56 188.66
DeFT-Flatten 9.98 10.99 12.48 94.67 66.95 10.90 16.10 42.23 46.60 56.96 84.27
Attention Speedup over Radix Attention 1.73$\times$ 1.63$\times$ 1.70$\times$ 1.39$\times$ 1.15$\times$ 1.21$\times$ 1.34$\times$ 1.96$\times$ 2.41×2.41$\times$2.41 × 3.11×3.$1$ 3.59×3.59$\times$3.59 ×
Decoding Speedup over Radix Attention 1.24$\times$ 1.28$\times$ 1.33$\times$ 1.10$\times$ 1.03$\times$ 1.03$\times$ 1.05$\times$ 1.29$\times$ 1.50×1.50$\times$1.50 × 1.91×1.91$\times$1.91 × 2.23×2.23$\times$2.23 ×
Speedup Upper-bound(no attention) 1.71$\times$ 2.08$\times$ 2.51$\times$ 1.96$\times$ 1.82$\times$ 1.70$\times$ 1.76$\times$ 1.89$\times$ 2.89$\times$ 3.34$\times$ 4.36$\times$
@@ 4.2 Analysis of Memory Management and Bottleneck
Figure 4: Latency breakdown for speculative decoding with a token tree of 32 queries, whose tree topology is from Medusa (Cai et al., 2024). U means unpaged memory.
@@
4.3 Decoding Behaviors: Latency and IOs.
We evaluate DeFT’s performance on various tree-based decoding tasks by measuring decoding latency (Table 5), which demonstrates DeFT’s acceleration of tree attention and wall clock time. See attention latency (Table 16), IO (Table 17), and inference accuracy (Table 15) in Appendix A.7. Notably, decoding latency essentially represents the optimal end-to-end (e2e) latency, excluding other overheads such as prefill latency (which accounts for approximately 5-10% of e2e latency) and framework-induced overheads (roughly 10-15% of e2e latency), including tree search, branching, etc. We exclude these overheads because they are consistent across all baselines and to eliminate the influence of the framework.
For few-shot prompting tasks, we used a prompt with 4k tokens and performed 400 decoding iterations, achieving a 1.33$\times$ decoding speedup thanks to 1.70$\times$ faster attention calculation and an approximately 90% reduction in IO.
For speculative decoding tasks, DeFT-Flatten achieved up to a 2.23$\times$ decoding speedup due to up to a 3.59$\times$ attention speedup, as the entire token tree (all queries) can share IO of the long prefix.
For multi-step reasoning tasks. While DeFT-Flatten improves attention speed by up to 1.36×1.36$\times$1.36 ×, the decoding acceleration is less pronounced due to two reasons:(1) The narrow tree width (10) restricts KV cache reuse, though increasing it to 50 in few-shot prompting significantly boosts decoding speed (1.2×1.2$\times$1.2 ×-1.5×1.5$\times$1.5 × over 100 iterations, see Appendix A.7). (2)The small number of tokens in the tree keeps attention at  30% of decoding latency, compared to 50-80% in speculative decoding. A longer prompt length increases attention computation overhead, leading to greater speedup, as shown in Table 8.
@@ 4.4 Ablation Study
Table 6: [Different KV Splitting Strategies] Comparison of DeFT-Node, DeFT-Node-Chunk and DeFT-Flatten in average attention latency (second) with NVIDIA A100 (80GB) for Llama3-8B model(GQA). This table is supplementary to Table 16. The fastest method is in bold, and the second fastest is underlined. Radix Attention is the best baseline in decoding latency. See details of more baselines in Table 16.
@@
Paged Radix Attention 5.99 7.30 9.96 39.37 24.69 3.11 5.13 25.73 40.47 76.10 145.43
DeFT-Node 10.59 10.62 10.85 42.96 33.29 6.16 9.58 34.59 34.41 34.96 41.78
DeFT-Node-Chunk 8.52 9.69 13.45 49.63 36.37 4.77 7.40 14.54 20.28 32.57 57.26
DeFT-Flatten 3.47 4.07 5.87 28.41 21.45 2.57 3.83 13.15 16.79 24.46 40.56
@@
See Table 8. With a longer prompt, DeFT-Flatten shows a more pronounced speedup in the same model, since the attention overhead is proportional to the token count in the decoding tree, while the FFN overhead remains nearly constant for the same model. See Figure 16, Figure 17 and Figure 18 for more results.
@@
See Table 8. With the larger model, Codellama-34B, DeFT-Flatten achieves slightly reduced but significant (up to 1.78$\times$) decoding speedup. The performance reduction is attributed to a lower A/F-LR, as the FFN overhead is greater due to larger hidden dimensions.
@@
Decoding 1.09$\times$ 1.37$\times$ 1.53$\times$ 1.67$\times$
Table 8: [Different Model Sizes] Comparison of decoding latency speedup and Attention/FFN latency ratio (in short as A/F-LR) between DeFT and Radix Attention for Codellama-34B and Codellama-7B models. Radix Attention is the best baseline in decoding latency. $b$ represents the tree width, and $t$ denotes the token tree size. For multi-step reasoning, we test the task sorting whose prompt length is about 1k tokens.
@@
Decoding Time Speedup 7B 1.34$\times$ 1.09$\times$ 1.85$\times$
@@ 5 Conclusion
We propose DeFT-Flatten, a hardware-efficient attention algorithm optimized for tree-structured LLM inference. It effectively addresses memory access and GPU utilization bottlenecks by reusing shared prefixes’ KV cache and evenly distributing workload across partitions. DeFT-Flatten’s key strengths lie in its prefix-sharing awareness and load balancing, making it versatile for various tree-structured tasks. It also scales well with larger search spaces and multiple branches. Our results show that DeFT-Flatten achieves up to 2.23×2.23$\times$2.23 ×/3.59×3.59$\times$3.59 × speedup in decoding and attention latency, outperforming baselines in tasks such as few-shot prompting, multi-step reasoning, and speculative decoding. Our ablation studies highlight that: (1) balanced partitioning is critical, (2) DeFT-Flatten delivers significant speedups across various LLM models and GPU architectures, and (3) DeFT-Flatten provides even greater speedups with larger token sizes (e.g., longer prompts) and more branches in tree-structured requests.
Acknowledgement
This work was supported in part by the National Science and Technology Major Project (No. 2022ZD0115101), Research Center for Industries of the Future (RCIF) at Westlake University, Westlake Education Foundation, and Westlake University Center for High-performance Computing.
@@ A.2 Discussion of Tree-based Decoding
The existing inference frameworks (Zheng et al., 2023; Gim et al., 2023) focused on tree-based decoding efficiency primarily aim to: (1) reduce memory footprints (Zheng et al., 2023) to enable larger batch sizes for higher throughput; (2) reuse the prompt cache (Gim et al., 2023) to avoid recomputation of the KV cache for faster time-to-first-token (TTFT). However, their designs do not specifically target reducing the latency of the entire decoding process. We observe that the tree-structured feature of LLM inference could provide us with some advantages to speed up the decoding itself.